HUBBLE SPACE TELESCOPE DAILY REPORT #1633 PERIOD COVERED: 0000Z (UTC) 05/13/96 - 0000Z (UTC) 05/14/96 Daily Status Report as of 135/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 6274 (Linking the HIPPARCOS to the Extragalactic = Referance Frame) The FGS#3 was used to make astrometeric observations of = extra-galactic object NEWEGOD-57 and several other reference stars in = order to help link the HIPPARCOS satellite reference frame to that of the = HST. The observations were completed as planned, and no anomalies were = reported. 1.2 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.3 Completed Four Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.4 Completed FOS/BL 6781 (Soft X-Rays and the UV Spectra of a = Complete QSO Sample) The Faint Object Spectrograph (Blue Side) was used to obtain = ultraviolet spectra of the low-Z quasi-stellar object PG1114+445. The = ionization continuum of this object has been studied using the ROSAT = satellite. The strengths and line profiles of strong and weak = ultraviolet emission and absorption lines and the ultraviolet continuum = were studied. The observations were completed as planned and no = anomalies were reported. 1.5 Completed Two Sets of WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: = Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.6 Completed FOS/RD 6059 (Imaging Low-Mass Companions of Nearby = Stars) The Faint Object Spectrograph (Red Side) was used to make = observations of low-mass nearby star Wolf359 to search for possible faint = companions. The bright star was placed behind an occulting bar. The = observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of WF/PC-2 5840 (Imaging Polarimetry of the Eta = and AG Car Nebulae) The WF/PC-2 was used to obtain polarimetric observations of the = nebula surrounding the luminous blue star HD93308 (aka, Eta Car Nebula). = This nebula was formed during massive outbursts that rivaled supernova = explosions in terms of its luminosity. The observations were executed as = scheduled, and no anomalies were noted 1.8 Completed Two Sets of WF/PC-2 6365 (A Snapshot Survey For = Companions and Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to obtain snapshot observations of Wolf-Rayet = stars WR39 and WR48 in order to look for companion stars and clusters. = Photometry was performed on the Wolf-Rayet stars and associated stars. = The observations were completed as planned, and no anomalies were = reported. 1.9 Completed FOS/RD/BL 6017 (Soft X-Ray Transients, Outburst in Black = Hole Accretion Disks) The Faint Object Spectrograph (Both the Blue and Red Sides) were = used to make target of opportunity observations of the cataclysmic = contact-binary star J1655-40 which is known to show soft X-ray = transients. These outburst are believed to be caused by the accretion = disk surrounding a black hole. The observations were completed as = scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Operations Notes: As reported in 1.9 above, the target of opportunity observations = have begun. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH